OpenStudy (anonymous):

Square root of (x+deltaX)+ 2 - square root of (x-2)/ delta x

OpenStudy (anonymous):

Like this?\[\sqrt{x+\Delta x}+2-\sqrt{\frac{x-2}{\Delta x}}\]

OpenStudy (anonymous):

Or like this? \[\sqrt{x+\Delta x}+2-\frac{\sqrt{x-2}}{\Delta x}\]

OpenStudy (anonymous):

Also, what is the question?

OpenStudy (anonymous):

Oh its like this \[\sqrt{(x+\Delta x)+2} -\sqrt{x+2} / \Delta x \]

OpenStudy (anonymous):

the delta x is over bother square roots.

OpenStudy (anonymous):

question is to find the limit as (delta) x approaches 0

OpenStudy (anonymous):

\[\frac{\sqrt{x+\Delta x+2}-\sqrt{x-2}}{\Delta x}\]

OpenStudy (anonymous):

yes but uh its x+2

OpenStudy (anonymous):

\frac{sqrt{x+Delta x}+2-sqrt{x+2}}{Delta x}\[\frac{\sqrt{x+\Delta x}+2-\sqrt{x+2}}{\Delta x}\]

OpenStudy (anonymous):

yes